# Can LLM Watermarks Robustly Prevent Unauthorized Knowledge Distillation?

### v1

@@ Abstract
The radioactive nature of Large Language Model (LLM) watermarking enables the detection of watermarks inherited by student models when trained on the outputs of watermarked teacher models, making it a promising tool for preventing unauthorized knowledge distillation. However, the robustness of watermark radioactivity against adversarial actors remains largely unexplored. In this paper, we investigate whether student models can acquire the capabilities of teacher models through knowledge distillation while avoiding watermark inheritance. We propose two categories of watermark removal approaches: pre-distillation removal through untargeted and targeted training data paraphrasing (UP and TP), and post-distillation removal through inference-time watermark neutralization (WN). Extensive experiments across multiple model pairs, watermarking schemes and hyper-parameter settings demonstrate that both TP and WN thoroughly eliminate inherited watermarks, with WN achieving this while maintaining knowledge transfer efficiency and low computational overhead. Given the ongoing deployment of watermarking techniques in production LLMs, these findings emphasize the urgent need for more robust defense strategies. Our code is available at https://github.com/THU-BPM/Watermark-Radioactivity-Attack.
@@
Leyi Pan1thanks: Work was done during the intern at Zhipu AI.,     Aiwei Liu1,     Shiyu Huang2,     Yijian Lu1,     Xuming Hu3,     Lijie Wen1thanks: Corresponding author,     Irwin King4,     Philip S. Yu5 1Tsinghua University    2Zhipu AI 3The Hong Kong University of Science and Technology (Guangzhou) 4The Chinese University of Hong Kong    5University of Illinois at Chicago panly24@mails.tsinghua.edu.cn, liuaw20@mails.tsinghua.edu.cn, wenlj@tsinghua.edu.cn
@@ 1 Introduction
Watermarking has emerged as a solution to monitor unauthorized usage Kirchenbauer et al. (2023); Zhao et al. (2024); Liu et al. (2024a); Zhao et al. (2023). Research has shown that watermarked LLMs exhibit radioactivity - student models trained on their outputs inherit detectable watermarks Sander et al. (2024); Gu et al. (2024). This traceability has led to increasing practical implementations, such as Google DeepMind’s integration of SynthID-Text Dathathri et al. (2024) into Gemini chatbots Team et al. (2023).
@@ 2.1 LLM Watermarking Schemes
Most of the existing watermarking schemes follow the $n$-gram paradigm, modifying the next token’s probability prediction based on the preceding $n-1$ tokens, thereby influencing the final sampling outcome Liu et al. (2024b); Kirchenbauer et al. (2023); Zhao et al. (2024); Dathathri et al. (2024); Liu et al. (2024a, 2024); Lee et al. (2024); Hu et al. (2024); Wu et al. (2023); Aaronson and Kirchner (2022); Kuditipudi et al. (2024); Pan et al. (2024); Liu et al. (2024); Lu et al. (2024). Watermark schemes tested in this work are:
@@ 2.2 Watermark Radioactivity
Research shows that watermarked LLMs exhibit radioactivity Sander et al. (2024); Gu et al. (2024): student models trained on their outputs inherit the watermark patterns. This effect is highly significant, with reported p-values below $10^{-30}$ even under the most stringent conditions where the teacher model is closed-source and detection is unsupervised (test prompts are disjoint from training data).
@@ 2.3 Watermark Removal Approaches
Prior work has explored various watermark removal methods, but focused on removing watermarks from generated text rather than models. These include untargeted approaches like paraphrasing, emoji attacks Kirchenbauer et al. (2023), back-translation, and cross-lingual removal He et al. (2024). For targeted removal, Jovanović et al. ; Wu and Chandrasekaran (2024); Zhang et al. (2024) proposed watermark stealing-and-removing, requiring knowledge of the type of watermarking scheme and the window size used.
@@ 3.1 Threat Model
Here are the threat model scenarios in which the student model holder acts as the attacker: (1) Attack Target: $n$-gram based watermarking schemes, which represent the primary approach in LLM watermarking (see Appendix D.2 for further discussion of other paradigms); (2) Access Level: Requires only access to the LLM service API, with no need for logits or watermark detection systems; (3) Testing Environment: Unsupervised testing, where test prompts differ from training data, as the LLM service cannot track which data was used to train the student model.
@@ 3.5.1 Watermark Radioactivity Factors
Scope of Watermark Stealing  Based on the preceding analysis, when conducting watermark stealing, we need only focus on scenarios with small values of $n$ (i.e., $n\leq 3$). Furthermore, for cases where $n\neq 1$, we can restrict our attention to high-frequency prefixes (i.e. those with frequencies exceeding $5\times 10^{-5}$).
@@ 4.1 Setup
Watermarking Schemes  KGW Kirchenbauer et al. (2023) and SynthID-Text Dathathri et al. (2024) with $n=1,2,3$. Results for more watermarking schemes can be found in Appendix D.1. All watermarking schemes were implemented using MarkLLM Pan et al. (2024), an open-source toolkit for LLM watermarking.
@@ 5.1 Computational Overhead
External Tools  UP and TP require external paraphrasing tools, while WN needs none.
Preprocessing Time Consumption  UP requires Dipper-like models to paraphrase the entire training dataset, WN only needs watermark stealing for preprocessing, and TP requires both. The time cost of watermark stealing consists of student model training (about 1h for Llama-7b) and dataset forward passes. The forward passes are computationally much cheaper than paraphrasing, as parallel computation is faster than autoregressive generation.
Student Model Inference Latency  UP and TP have no additional inference overhead. WN requires adding inverse watermark during inference, causing slight delay.
@@ 5.3 Future Directions in Defense Strategy
This work reveals the vulnerability of unauthorized knowledge distillation prevention when generative LLM watermarking is predominantly confined to n-gram based approaches. While alternatives such as sentence-level reject sampling Hou et al. (2023, 2024) and post-generation signal embedding Chang et al. (2024) exist, these approaches introduce significant latency to the current token-by-token real-time LLM inference paradigm, making them difficult to deploy at scale in real-time LLM services. We therefore advocate for diversifying token-level watermarking techniques, as multiple paradigms would make it harder for attackers to identify and target specific methods, enabling more robust protection.
@@ 6 Conclusion
This work presents the first systematic study of the robustness of watermarking schemes against adversarial attacks in preventing unauthorized knowledge distillation. We propose three watermark removal approaches: two pre-distillation methods (UP, TP) and one post-distillation method (WN). Through comprehensive experiments, we evaluate the resilience of watermarking schemes against these attacks. Our findings reveal that WN achieves effective watermark removal while maintaining superior knowledge preservation, highlighting the urgent need for more robust defensive strategies.
@@ Appendix F Adaptive Control for Inverse Watermark Strength
However, to account for potential extreme cases, we also explored strategies for adaptive control of inverse watermark strength. Our approach is to estimate the required inverse watermark strength $\delta^{\prime}$ by detecting the watermark intensity inherited by the student model. Since the student model holder does not have access to the watermark detector, we employed Water-Probe Liu et al. (2025) to measure watermark intensity. Water-Probe is a recently proposed identification algorithm that tests for watermarks by comparing the model’s responses to specially crafted prompts, where higher similarity in responses to crafted prompts pairs indicates a higher likelihood (or strength) of watermarking.

### v2

@@ Abstract
The radioactive nature of Large Language Model (LLM) watermarking enables the detection of watermarks inherited by student models when trained on the outputs of watermarked teacher models, making it a promising tool for preventing unauthorized knowledge distillation. However, the robustness of watermark radioactivity against adversarial actors remains largely unexplored. In this paper, we investigate whether student models can acquire the capabilities of teacher models through knowledge distillation while avoiding watermark inheritance. We propose two categories of watermark removal approaches: pre-distillation removal through untargeted and targeted training data paraphrasing (UP and TP), and post-distillation removal through inference-time watermark neutralization (WN). Extensive experiments across multiple model pairs, watermarking schemes and hyper-parameter settings demonstrate that both TP and WN thoroughly eliminate inherited watermarks, with WN achieving this while maintaining knowledge transfer efficiency and low computational overhead. Given the ongoing deployment of watermarking techniques in production LLMs, these findings emphasize the urgent need for more robust defense strategies.
@@
additions
Leyi Pan1, Aiwei Liu1†, Shiyu Huang2, Yijian Lu1, Xuming Hu3, Lijie Wen1†, Irwin King4, Philip S. Yu5 1Tsinghua University, 2Zhipu AI, 3The Hong Kong University of Science and Technology (Guangzhou), 4The Chinese University of Hong Kong, 5University of Illinois at Chicago panly24@mails.tsinghua.edu.cn, liuaw20@mails.tsinghua.edu.cn, wenlj@tsinghua.edu.cn
footnotetext: Corresponding authors.
@@ 1 Introduction
Watermarking has emerged as a solution to monitor unauthorized usage Kirchenbauer et al. (2023); Zhao et al. (2024); Liu et al. (2024c); Zhao et al. (2023). Research has shown that watermarked LLMs exhibit radioactivity - student models trained on their outputs inherit detectable watermarks Sander et al. (2024); Gu et al. (2024). This traceability has led to increasing practical implementations, such as Google DeepMind’s integration of SynthID-Text Dathathri et al. (2024) into Gemini chatbots Team et al. (2023).
@@ 2.1 LLM Watermarking Schemes
Most of the existing watermarking schemes follow the $n$-gram paradigm, modifying the next token’s probability prediction based on the preceding $n-1$ tokens, thereby influencing the final sampling outcome Kirchenbauer et al. (2023); Zhao et al. (2024); Dathathri et al. (2024); Liu et al. (2024c, b); Lee et al. (2024); Hu et al. (2024); Wu et al. (2023); Aaronson and Kirchner (2022); Kuditipudi et al. (2024); Liu et al. (2024d); Lu et al. (2024); Pan et al. (2025). Watermark schemes tested in this work are:
@@ 2.2 Watermark Radioactivity
Research has shown that watermarked LLMs exhibit "radioactivity" — their distinctive watermark patterns can be inherited by student models trained on their outputs Sander et al. (2024); Gu et al. (2024). This characteristic is especially valuable for detecting unauthorized knowledge distillation. If a more capable teacher model employs watermark, any training data collected from its API will inherently carry the watermark. Consequently, student models trained on this data will produce outputs that retain the same watermark patterns, allowing the teacher model’s owner to detect and trace unauthorized knowledge distillation attempts.
The underlying reason why “the watermark has radioactivity" is that the watermark itself introduces a pattern that biases certain subsequent tokens based on the n-grams of the preceding context. When a student model is trained on watermarked data, it adopts this same bias. This effect is remarkably significant, with reported p-values below $10^{-30}$ even under the most stringent conditions: (1) the teacher model is closed-source, meaning student models can only learn from the output text and cannot access the teacher model’s logits, which would otherwise make it easier to learn the pattern; and (2) watermark detection is performed unsupervised, meaning the test prompts for the student models are entirely disjoint from the training data.
@@ 2.3 Watermark Removal Approaches
Previous research has explored diverse approaches to watermark removal, primarily concentrating on eliminating watermarks from generated text rather than from the models themselves. These approaches can be categorized into untargeted and targeted methods. The untargeted approaches encompass various text transformation techniques, including paraphrasing, emoji-based attacks Kirchenbauer et al. (2023), back-translation, and cross-lingual removal strategies He et al. (2024). In the realm of targeted removal, researchers such as Jovanović et al. ; Wu and Chandrasekaran (2024) and Zhang et al. (2024) have proposed watermark stealing-and-removing techniques, though these methods necessitate prior knowledge of both the specific watermarking scheme employed and its corresponding window size parameter.
@@ 3.1 Threat Model
Overview of the Attack Scenario  Assume there is a closed-source teacher model, which employs a watermarking scheme, making all outputs acquired from its API carries the watermark. Now, consider the owner of a student model, who seeks to perform knowledge distillation by training on teacher’s watermarked outputs. The owner applies various watermark removal attacks—such as untargeted training data paraphrasing, targeted training data paraphrasing, and watermark neutralization—with the goal of removing the inherited watermark from the trained student model, while preserving the knowledge it has learned.
Watermarking Schemes Examined  This investigation encompasses all n-gram-based watermarking schemes, including KGW Kirchenbauer et al. (2023), SynthID-Text Dathathri et al. (2024), KGW-Minhash Kirchenbauer et al. (2024), KGW-SkipHash Kirchenbauer et al. (2024), Unbiased WatermarkHu et al. (2024), DiPMark Wu et al. (2023), Aar Aaronson and Kirchner (2022), and SIR Liu et al. (2024c), among others. These schemes represent the predominant paradigm of contemporary LLM watermarking methodologies. In our primary experiments, we have presented experimental findings for KGW and SynthID-Text. Furthermore, Appendix D contains supplementary results pertaining to KGW-Minhash and KGW-SkipHash.
Access Requirements for the Student Model Owner to Perform Watermark Removal The student model owner requires the following: (1) Access to the training data (2) Access to both the original and trained student models. Since our removal methods are designed from the student model’s perspective, these requirements are practical and feasible. Importantly, the student model owner does NOT need access to the watermark detection system or its API, which are controlled by the teacher model owner.
Watermark Detection The watermark detection for the trained student model is performed by the owner of teacher model, involving prompting the trained student model to generate a certain amount of output text. These outputs are then analyzed to detect the presence of the watermark and to compute a confidence score. The test prompts are distinct from the training data, assuming that the teacher model service cannot track which specific data was used to train the student model.
@@ 3.5.1 Watermark Radioactivity Factors
Scope of Watermark Stealing  Based on the preceding analysis, when conducting watermark stealing, we need only focus on scenarios with small values of $n$ (i.e., $n\leq 3$). That is to say, prior knowledge of the window size employed in the teacher model’s watermarking algorithm is not required; instead, we can directly iterate through a small range of $n$ values, as detailed in Section 3.5.2. Furthermore, for cases where $n\neq 1$, we can restrict our attention to high-frequency prefixes (i.e. those with frequencies exceeding $5\times 10^{-5}$).
@@ 4.1 Setup
Watermarking Schemes  KGW Kirchenbauer et al. (2023) and SynthID-Text Dathathri et al. (2024) with $n=1,2,3$. Results for more watermarking schemes can be found in Appendix D. All watermarking algorithms tested are implemented using the MarkLLM toolkit Pan et al. (2024).
@@ 5.1 Computational Overhead
External Tools  Both UP (Untargeted Paraphrasing) and TP (Targeted Paraphrasing) require external paraphrasing tools for text transformation, which introduces additional dependencies and potential costs. In contrast, WN (Watermark Neutralization) operates independently without the need for external tools, making it more self-contained and easier to deploy.
Preprocessing Time Consumption  The preprocessing requirements vary significantly among the three methods. UP necessitates the use of Dipper-like models to paraphrase the entire training dataset, which can be computationally intensive and time-consuming. WN’s preprocessing is more efficient, requiring only watermark stealing, which involves student model training (approximately 1 hour for Llama-7b) and dataset forward passes. TP combines both requirements, making it the most time-intensive in preprocessing. Notably, the forward passes required for watermark stealing are computationally more efficient than paraphrasing operations, as they can leverage parallel computation rather than relying on slower autoregressive generation.
Student Model Inference Latency  From an inference perspective, UP and TP maintain the original model’s performance with no additional overhead during inference. WN introduces a minor latency due to the necessity of adding inverse watermark during inference. However, this additional computational cost is relatively minimal and may be acceptable in many practical applications.
@@ 5.3 Future Directions in Defense Strategy
This work reveals the vulnerability of unauthorized knowledge distillation prevention when generative LLM watermarking is predominantly confined to n-gram based approaches. While alternatives such as sentence-level reject sampling Hou et al. (2023, 2024) and post-generation signal embedding Chang et al. (2024) exist, these approaches introduce significant latency to the current token-by-token real-time LLM inference paradigm, making them difficult to deploy at scale in real-time LLM services.
Given these constraints, we advocate for a more diverse ecosystem of token-level watermarking techniques. The development of multiple paradigms would create a more complex landscape for attackers to navigate, while maintaining the efficiency required for real-world applications. Future research should focus on novel token-level approaches that can be seamlessly integrated into existing LLM inference pipelines while providing robust protection against various forms of attacks.
@@ 6 Conclusion
This work presents the first systematic study of the robustness of watermarking schemes against adversarial attacks in preventing unauthorized knowledge distillation. We propose three watermark removal approaches: two pre-distillation methods (Untargeted Paraphrasing, Targeted Paraphrasing) and one post-distillation method (Watermark Neutralization). Through comprehensive experiments, we evaluate the resilience of watermarking schemes against these attacks. Our findings reveal that WN achieves effective watermark removal while maintaining superior knowledge preservation, highlighting the urgent need for more robust defensive strategies.
@@
Acknowledgments
This work is primarily supported by the Key Research and Development Program of China (No. 2024YFB3309702). We would like to express our gratitude to the anonymous ARR Feburary reviewers (Reviewer mdgr, LhPY, 7gDE) and Area Chair LGVF for their valuable feedback and suggestions that helped improve this paper.
@@ D.2 Discussion About Other Watermarking Paradigms
D.3 Discussion About Repeated Context Masking
Another noteworthy watermarking technique to examine is repeated context masking, where watermarks are only applied to subsequent tokens when a context is first encountered, but not when similar contexts appear later in the sequence. This approach has been implemented in several watermarking algorithms, including SynthID-Text Dathathri et al. (2024) (which offers both a distortionary version and a non-distortionary version utilizing repeated context masking, with our main experiments employing the distortionary variant) and Unbiased watermark Hu et al. (2024). The technique is specifically designed to ensure the watermarking process remains distortion-free.
Using repeated context masking could significantly reduce watermark radioactivity. As demonstrated in Figure 3, watermark radioactivity exhibits a strong correlation with prefix frequency in the training data. When repeated context masking is implemented, the frequency of watermark patterns in the training data diminishes substantially, thereby impeding the student models’ ability to acquire these patterns.
We conducted experimental evaluations of both SynthID-Text and Unbiased Watermark employing repeated context masking, while maintaining other parameters consistent with the main experiments. In the implementation of repeated context masking, we established a maximum storage capacity of 1024 distinct contexts. For student models without any watermark removal attack, the median p-value of watermark detection are shown in Table 11.
Table 11: Median p-value of watermark detection in student models trained on texts watermarked with repeated context masking.
Watermarking Scheme Token Num. Llama-7b Llama-3.2-1b
SynthID-Text (w. Context Mask) $n=2$ 10k 1.50e-01 1.10e-01
20k 1.00e-01 8.34e-02
30k 3.00e-02 2.11e-02
$n=3$ 100k 3.50e-01 2.80e-01
300k 2.80e-01 2.20e-01
1 million 2.10e-01 1.70e-01
UW (w. Context Mask) $n=2$ 10k 2.10e-01 1.80e-01
20k 1.60e-01 1.30e-01
30k 9.50e-02 7.20e-02
$n=3$ 100k 4.50e-01 3.80e-01
300k 3.70e-01 3.00e-01
1 million 3.00e-01 2.50e-01
@@ Appendix F Adaptive Control for Inverse Watermark Strength
However, to account for potential extreme cases, we also explored strategies for adaptive control of inverse watermark strength. Our approach is to estimate the required inverse watermark strength $\delta^{\prime}$ by detecting the watermark intensity inherited by the student model. Since the student model holder does not have access to the watermark detector, we employed Water-Probe Liu et al. (2024a) to measure watermark intensity. Water-Probe is a recently proposed identification algorithm that tests for watermarks by comparing the model’s responses to specially crafted prompts, where higher similarity in responses to crafted prompts pairs indicates a higher likelihood (or strength) of watermarking.